OpenStudy (anonymous):

A sandbag was thrown downward from a building. The function f(t) = -16t2 - 32t + 128 shows the height f(t), in feet, of the sandbag after t seconds:

OpenStudy (anonymous):

Part A: Factor the function f(t) and use the factors to interpret the meaning of the x-intercept of the function. Part B: Complete the square of the expression for f(x) to determine the vertex of the graph of f(x). Would this be a maximum or minimum on the graph? Part C: Use your answer in part B to determine the axis of symmetry for f(x)?

OpenStudy (anonymous):

Do you know how to factor?

OpenStudy (anonymous):

a little bit

OpenStudy (anonymous):

@ganeshie8 @tHe_FiZiCx99 @jim_thompson5910 @bahrom7893 @iambatman @Ashleyisakitty

OpenStudy (ranga):

-16t^2 - 32t + 128 Factor out -16 first. What do you get?

OpenStudy (anonymous):

I don't know

OpenStudy (ranga):

Take -16 out and divide each coefficient by -16

OpenStudy (anonymous):

2?

OpenStudy (ranga):

There are 3 coefficients: -16, -32, +128 When you factor out -16 you have to divide each coefficient by -16.

OpenStudy (anonymous):

oooohh okay

OpenStudy (anonymous):

@ranga now what do I do?

OpenStudy (ranga):

Factor out -16 from -16t^2 - 32t + 128 by dividing each coefficient by -16: -16t^2 - 32t + 128 = -16( ? t^2 + ? t + ?) fill in the 3 question marks by dividing the corresponding coefficient by -16

OpenStudy (anonymous):

what do I do now?*

OpenStudy (ranga):

You should know how to factor out a constant by now. You may want to brush up because you are likely to come across factoring again and again.

OpenStudy (anonymous):

−16(t+4)(t−2)

OpenStudy (ranga):

Yes.

OpenStudy (ranga):

x-intercept is when f(t) = 0 -16(t+4)(t-2) = 0 t = ?

OpenStudy (anonymous):

idk :/

OpenStudy (ranga):

If (x+a)(x+b) = 0 it implies x + a = 0 or x + b = 0 x = -a or x = -b

OpenStudy (anonymous):

ok..

OpenStudy (ranga):

-16(t+4)(t-2) = 0 when t = ?

OpenStudy (anonymous):

-4?

OpenStudy (anonymous):

-6?

OpenStudy (ranga):

-16(t+4)(t-2) = 0 implies t + 4 = 0 OR t - 2 = 0 t = -4 OR t = 2 Here t is a measure of time in seconds. So negative time does not make sense and can be ignored. t = 2 seconds is the solution. "interpret the meaning of the x-intercept of the function" x-intercept is when y or f(t) is zero. Therefore, when t = 2, f(t) = 0. f(t) is the height of the sandbag at time t. f(t) = 0 implies the height is zero or the sandbag hits the ground. So at t = 2 seconds, the sandbag reaches the ground.

OpenStudy (anonymous):

ok

OpenStudy (ranga):

f(t) = -16t^2 - 32t + 128 = −16(t+4)(t−2) = -16( t^2 + 2t - 8) complete the square of ( t^2 + 2t - 8)

OpenStudy (anonymous):

wait what????

OpenStudy (ranga):

In part A I ask you to factor -16 out of -16t^2 - 32t + 128. -16t^2 - 32t + 128 = -16(t^2 + 2t - 8) In part B they are asking you to complete the square. So complete the square of ( t^2 + 2t - 8)

OpenStudy (anonymous):

i don't remember the complete the square method

OpenStudy (ranga):

take the coefficient of the middle term and divide by 2 t^2 + 2t - 8. Coefficient of middle term is +2. Divide by 2 to get 1. This 1 will go inside the parenthesis to be squared and outside you have to subtract 1^2 t^2 + 2t - 8 = (t + 1)^2 - 1^2 - 8 = (t + 1)^2 - 9 f(t) = -16t^2 - 32t + 128 = -16(t^2 + 2t - 8) = -16{ (t + 1)^2 - 9 } = f(t) = -16(t+1)^2 + 144 compare this to the vertex for of a parabola y = a(x-h)^2 + k where (h,k) is the vertex and you get h = -1 and k = 144. So the vertex is at (-1, 144). The vertex will be a maximum because the leading coefficient of the parabola is -16 which is negative and therefore the parabola open downward which means the vertex is a maximum

OpenStudy (ranga):

Part C) This is a vertical parabola which means the axis of symmetry will be a VERTICAL line passing through the vertex. The equation of a vertical line passing through (-1, 144) is x = -1.

OpenStudy (anonymous):

thanks i appreciate the help